Obertext zum Ganzen

Text1  Text1a
\HL  horizontal line
\VL  vertical line
\NR  next row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row

Obertext zum Ganzen

\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row
\SR  single row
\FR  first row
\MR  middle row
\LR  last row

Table 1 sample table